RC meeting and Run Plan, Sunday February 19th
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Report
- https://logbooks.jlab.org/entry/4134791
- RC: Bogdan Wojtsekhowski, bogdanw@jlab.org
Summary
- Recovery from the fire event was successful
- Ran with ~45 uA beam overnight
- Owl shift on 2/19 got 1 C - which is very good
Issues
- DAQ recovery procedure is a big puzzle
- Coordination of the "experts on call" - e.g. TH needs improvement
- Beam trip rate is high and beam position moves a lot
Current Status
- Production on He-3 target
Run Plan
- Continue running 45 uA on He3
- Accelerator will prioritize beam and not have any maintenance days this week
Data Collection and Polarization Status
- 37.1C of uncorrected charge for this kinematic (42.8C of corrected charge which is 27% of the plan)
- Last NMR 49.3%
Other Topics/Status Updates
- Anything else?
- Bill wants to change the raster settings after some analysis
Requested Accesses to the Hall
- Jack would like to collect some cables